find the residue of the function f(z)=1/z(z-2) where c is the circle |z|=1


1
Expert's answer
2021-09-01T11:56:32-0400

"f(z)=\frac{1}{z(z-2)}=\frac{1}{2}(\frac{1}{z-2}-\frac{1}{z})"

The function "\frac{1}{2}\frac{1}{z-2}" has a unique pole at "z=2" and has no singularities inside the contour "C=\{z\in\mathbb{C}:|z|=1\}". Therefore, the residue of this function equal to 0.

The function "-\frac{1}{2z}" has a unique pole at "z=0" and it is inside the contour C. The coefficient with "z^{-1}" is equal to "-1\/2", therefore, the residue of this function is equal to "-1\/2".

Therefore, the residue of the function "f(z)=\frac{1}{z(z-2)}" inside the contour "C" is equal to "0-1\/2=-1\/2".

Answer. -1/2